Grade 11Trigonometry

solve for x
sin3x *sin(cube) x +cos3x *cos(cube) x=0

rashmi
12 Years agoGrade 11

2 Answers

Arun Kumar
12 Years ago
Hello Student,

sin3x *sin^3 x +cos3x *cos^3 x=0
tan3x*tan^3x=0
=>tan3x=0 union tan^3x=0
=>3x=n*pi union x=n*pi
rashmi
12 Years ago
please give some easy method as i am in class 11th this iss going above my head or can u just explain it